OpenStudy (anonymous):

Give and example ( by writing an equation) of each of the following: a function whose domain is [0, infinity) a function whose domain is ( -infinity, infinity) a function whose domain is (-infinity, 0) u ( 0, infinity)

OpenStudy (anonymous):

help

OpenStudy (anonymous):

@sweetburger @Gemini_Nation_ @Firez @fishes220 help someone

OpenStudy (anonymous):

help

OpenStudy (anonymous):

@IrishBoy123 some one help i give medals!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

ill help

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

k

OpenStudy (anonymous):

f(x) = x + 1 for number 2

OpenStudy (anonymous):

f(x) = x + 1 for number 2

OpenStudy (anonymous):

1-x if x is greater than or equal to zero. third 1

OpenStudy (anonymous):

and finally the first 1 is f(x)= x +4 if x is less than zero

OpenStudy (anonymous):

did this help @Julianne6th

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is this right

OpenStudy (anonymous):

You can't take the square root of negative numbers, so the domain of \[f(x) = \sqrt x\] is: \[[0,\infty)\]

OpenStudy (anonymous):

What is the domain of this function: \[f(x) = 1/x\] (Think: what number(s) are you not allowed to plug in for x?)

OpenStudy (anonymous):

this is @cookiimonster627 question i posted it for him or her

OpenStudy (anonymous):

just FYI: the domains of both f(x) = 1-x and f(x) = x + 4 are all real numbers: that is, the interval (-infinity,infinity). Why? What number(s) can be plugged into 1 - x ? Any numbers! That is, you can subtract *any* number from 1. (and so on)

OpenStudy (anonymous):

@OneMathCat your so wrong that shows your not a math cat